# Solving the Bongard-logo Problem by Modeling a Probabilistic Model

### v7

@@ IV-B Straw Pose-Tansformer
Pose-Transformer, during deployment, exerts considerable pressure on computational resources. The establishment and computation of the mapping matrices, denoted as $\{W_{ijk}|i\in[1,n],j\in[1,m],k\in[1,n],W_{ijk}\in R^{(d/m)\times d}\}$, incur significant resource overhead. Motivated by this challenge, this paper explores the lightweighting of Pose-Transformer. The crux of Pose-Transformer lies in computing compressed pose vectors and subsequently employing them to modify and refine the logical inputs, represented as $\{V_{i}|i\in[1,n],V_{i}\in R^{d}\}$ to some extent. The computation of the logical outputs in Pose-Transformer’s pose matrix mapping follows the methodology outlined below:
@@
It is worth noting that the placeholders marked with “_” in the formula represent numbering for on-demand broadcasting. Multiple levels of summation and mapping can incur significant computational overhead. By designing parameter sharing, we aim to eliminate partial logical summation and mapping, thereby achieving the purpose of reducing computational cost. From a logical hierarchy perspective, only the numbering at the $i$-th level has the potential for indentation, while other logical levels have strong mapping requirements. In other words, this paper posits that the pose matrices between equivalent local vectors across different tokens merit weight sharing. The planned target expression of pose matrix mapping after indentation is:
@@
This paper aims to simplify the Pose-Transformer based on the aforementioned expression. We treat the mapping matrix $W_{ijk}$ as a product of the mapping matrix $\tilde{W}_{jk}$ multiplied by $i$ different coefficients, specifically expressed as $W_{ijk}\tilde{=}w_{i}\cdot\tilde{W}_{\_jk}$. This transformation significantly reduces the number of matrices involved. We argue that it is more reasonable to combine $w_{i}$ with $H_{ij\_}$ and treat them as a weighted sum of $H_{ij\_}$, rather than treating $w_{i}$ as an optimizable weight vector. Furthermore, treating this process as computing an attention-weighted sum for $H_{ij\_}$ is even more elegant and effective. Therefore, we propose a feasible solution that involves embedding a learnable vector $V_{0}$, where $V_{0}\in R^{d}$ and we denote $\{H_{0j}|j\in[1,m],H_{0j}\in R^{(d/m)}\}$ as the attention heads corresponding to $V_{0}$, and computing the multi-head cross-attention results between $V_{0}$ and $\{V_{i}|i\in[1,n]\}$. This approach aligns well with the formula presented:
@@
Notably, the aforementioned approach necessitates computing cross-attention additionally between the standard Transformer-Encoder layer and pose mapping, which prolongs the training time, contradicting the intended purpose. This paper utilizes a transformer encoder with an additional mask mechanism to integrate the transformer encoding process with the multi-head cross attention computation process.
Specifically, this paper introduces an optimizable vector, denoted as $\tilde{V}_{0}\in R^{d}$, for the embedding of logical inputs $\{\tilde{V}_{i}|i\in[1,n]V_{i}\in R^{d}\}$. This vector is akin to the class learning vector in the Vision Transformer (ViT). After the logical inputs, augmented with a learnable vector $\tilde{V}_{0}$, propagate through a single Transformer-Encoder layer while applying a mask to the self-attention weights at the position corresponding to $\tilde{V}_{0}$ to prevent its value mapping from contributing to the attention results of the Transformer-Encoder, they produce an updated set of vectors $\{V_{i}|i\in[0,n],V_{i}\in R^{d}\}$ representing the attention-weighted outputs. Our aim is to accomplish two processes within a single self-attention computation: the extra multi-head cross-attention between $\tilde{V}_{0}$ and $\{\tilde{V}_{i}|\,i\in[1,n]\}$ , and the pre-existing multi-head self-attention among $\{\tilde{V}_{i}|\,i\in[1,n]\}$. We replace the tedious sequential process with this masked attention mechanism, thereby streamlining the computational workflow. Due to the inherent self-attention mechanism and masking, the vector $\tilde{V}_{0}$ assimilates information from the logical inputs, transforming into global vector $V_{0}$. Within this global vector, we establish a collection of mapping matrices $\{\tilde{W}_{jk}|j\in[1,m],k\in[1,n],\tilde{W}_{jk}\in R^{(d/m)\times d}\}$ for each attention head $\{H_{0j}|j\in[1,m],H_{0j}\in R^{(d/m)}\}$ in it. It is evident that there exists a subtle difference between the masked attention process and the sequential process of self-attention and cross-attention, which renders them not strictly equivalent. Conversely, the masked attention process is equivalent to a serial process of performing cross-attention followed by self-attention. However, ignoring the sequence and merging the two processes of attention-based weighted summation into one step is not an unreasonable approach. In fact, this merger can be seen as a computational optimization that reduces the amount of computation and improves efficiency.
@@
The pose vectors $\{p_{jk}|j\in[1,m],k\in[1,n],p_{jk}\in R^{d}\}$ is achieved by multiplying the attention heads with the mapping matrices. These pose vectors, after summation and squashing, are added to the vector set $\{V_{i}|i\in[1,n],V_{i}\in R^{d}\}$ to obtain the output of new pose matrix mapping block $\{O_{i}|i\in[1,n],O_{i}\in R^{d}\}$. This process outlines the computation process of the pose embedding matrix block within the Straw-Pose-Transformer architecture. The entire calculation process of new “pose matrix embedding” can be expressed as fallows:
@@
In the formula, the process of the Masked transformer-encoder to obtain the output is denoted as “Masked T-Encoder”. In the Straw-Pose-Transformer architecture, the outputs from each logical layer, denoted as $\{O_{k}|k\in[1,n]\}$, undergo a concatenation process with the global vector $V_{0}$ attached to their precursors. This modified set, now represented as $\{\tilde{V}_{i}|i\in[0,n]\}$ , is then propagated into the subsequent logical layer for further processing. However, if the subsequent logical layer is the final layer, the global vector $V_{0}$ is no longer appended.
In this way, we can reconstruct the pose matrix embedding process according to the formula, and this feedforward process is illustrated in the figure 10. It is worth mentioning that masking the position corresponding to $V_{0}$ in the attention coefficients reduces the impact of the optimizable parameter $V_{0}$ on the pose mapping process to a certain extent, making it a relatively necessary operation. our analysis demonstrates that this stabilizes training and improves inference accuracy somewhat. This paper replaces the standard Transformer-Encoder+cross-attention+pose matrix mapping sequential process in Pose-Transformer with the aforementioned procedure. And figure 11 illustrate the structure of the Straw-Pose-Transformer.
Figure 10: The feedforward process of a single Straw pose matrix embedding block
@@
In addressing the Bongard-Logo problem, this paper introduces a novel baseline, PMoC. To demonstrate the adaptability of the Straw-Pose-Transformer in solving the Bongard-Logo problem, the Transformer-Encoder component within PMoC is replaced with the Straw-Pose-Transformer in this section. Furthermore, given the remarkable performance achieved by the predecessor RS-Tran in solving the RPM problem, this paper does not design an additional baseline network to prove the superiority of the Pose-Transformer in addressing the RPM problem. This paper aims to demonstrate the advantage of Pose-Transformer over the standard Transformer-Encoder for RPM problems by achieving improved accuracy through the simple replacement of the Transformer-Encoder in RS-Tran with Pose-Transformer.
The RS-Tran[31] primarily consists of Transformer-Encoder structures, and in this study, we replace it with the Pose-Transformer and Straw-Pose-Transformer and conduct experiments. The modified RS-Tran is renamed as RS-Pose-Tran and RS-Straw-Pose-Tran. Due to the fundamental modifications made to the RS-Tran architecture, RS-Pose-Tran and RS-Straw-Pose-Tran can be regarded as models designed based on our novel structure, incorporating inductive biases inherent in the RPM problem[37]. Subsequently, it was discovered that compared to the Transformer-Encoder, the Pose-Transformer exhibits greater suitability for abstract reasoning problems, specifically the RPM and Bongard-Logo problems.
@@ VI-A Experiment on Bongard-Logo
We conducted experiments on the Bongard-logo dataset using the SBSD, PMoC and the integration of PMoC with Straw-Pose-Transformer. The results are recorded in Table I. Due to the serial convolution scanning mechanism and attention mechanism being prone to attention collapse, the parameters of the SBDC model are utilized as pre-trained parameters for the PMoC convolutional layer. It is worth noting that, under the experimental conditions of this paper, which involved four A100s graphics processing units (GPUs), the integration of PMoC with the Pose-Transformer proved to be challenging to deploy. This is one of the reasons and motivations behind our decision to design a lightweight approach for Pose-Transformer in this paper.
All our experiments are implemented in Python using the PyTorch[38] framework. In the experiments conducted for PMoC and SBSD, the Adam[39] optimizer was selected with a learning rate set to 0.001, a learning rate decay of 0.995, a weight decay of 0.0001, and a batch size of 40 for PMoC and SBSD. Specifically, in the experiment combining PMoC with Straw-Pose-Transformer, the batch size was set to 20.
@@
This paper argues that the adopted data augmentation methods do not explicitly alter the core reasoning difficulty or logical complexity of the Bongard-Logo problem. These augmentation techniques, from an intuitive perspective, neither add nor remove any preexisting concepts. Even if these data augmentation methods implicitly introduce additional and unknown concepts to the Bongard-Logo dataset, it is important to note that such concepts are not present in the test set, as no data augmentation has been applied to it. Furthermore, they do not provide guidance for generalization issues such as NV and CM, as the solutions to these problems are not directly reflected in the HD problem after applying such augmentations. Additionally, the data augmentation approach employed in this study is uniformly applied to all 14 images within each Bongard-Logo instance, maintaining the nature of few-shot learning (i.e., without increasing the number of images within an instance) and without introducing additional priors (such as rotational and flipping invariance of individual images within Bongard-Logo instances). These augmentation methods indeed enrich the expressions of concepts at the pixel level, which is precisely the challenge we anticipate for PMoC through data augmentation and expanded data volume.
@@ VI-B Experiment on RPM
We conducted experiments on I-RAVEN and PGM by sequentially replacing the Transformer structure in RS-Tran [31] with Pose-Transformer and Straw-Pose-Transformer. The results of I-RAVEN are recorded in Table III and PGM is recorded in Table IV. In this paper, we considered that compared to the standard Transformer-Encoder, a single layer of Pose-Transformer and its lightweight version, Straw-Pose-Transformer, possess twice the number of logical computations. Therefore, the combined RS-Pose-Tran and RS-Straw-Pose-Tran only stacked half of the logical layers of RS-Tran to ensure that the internal logical computation times of the three models were equivalent. In addition, the parameter settings and dataset configurations for the experiments related to RS-Pose-Tran were strictly maintained consistent with those of RS-Tran. Such a setting renders the comparative results of the three models more convincing.
@@
In Table IV, it can be seen that we did not conduct experiments on RS-Pose-Tran for PGM. This is because compared to RS-Tran operating on the RAVEN dataset, the overall dimension of the RS-Tran model doubles on PGM. However, such a doubling is prohibitive for RS-Pose-Tran, making it undeployable on our equipment and device. This also motivated us to lightweight the Pose-Transformer into the Straw-Pose-Transformer in this paper.
@@
From the perspective of experimental results, it is not difficult to find that our lightweight approach, RS-Straw-Pose-Tran, has been successful without significantly reducing the performance of RS-Pose-Tran. Furthermore, the introduction of the pose mapping matrix into the Transformer-Encoder is effective, significantly improving the performance of RS-Tran. This demonstrates that Pose-Transformer is more suitable for abstract reasoning capabilities compared to the standard Transformer-Encoder, even though the standard Transformer-Encoder structure has shown strong advantages in RPM and Bongard-logo problems.
@@ VII Conclusion
In this paper, we presented an empirical study on the Bongard-logo dataset using the PMoC model and conducted experiments on I-RAVEN and PGM by replacing the Transformer structure in RS-Tran with Pose-Transformer and Straw-Pose-Transformer, sequentially.
Our experimental results demonstrate that PMoC has achieved competitive results on the Bongard-logo problem, successfully tackling it in the form of a probabilistic model. The two compromises made in the model design for implementing the PMoC are effective and reasonable. And the results also show that the introduction of the pose mapping matrix into the Transformer-Encoder is effective, achieving a significant breakthrough in the already remarkable accuracy of RS-Tran. These results suggest that Pose-Transformer is more advantageous for abstract reasoning tasks compared to the standard Transformer-Encoder, even though the latter has exhibited strong capabilities in RPM and Bongard-logo problems. Furthermore, our lightweight approach has been successful without significantly reducing the performance of RS-Pose-Tran.
In conclusion, our Pose-Transformer emphasizes the effectiveness of integrating the pose mapping matrix into the Transformer-Encoder framework, thereby elevating abstract reasoning capabilities. Furthermore, we have successfully deployed a streamlined variant of Pose-Transformer, termed Straw-Pose-Transformer, without compromising performance, which further reduces resource consumption in addressing complex reasoning challenges. Additionally, the innovative design of PMoC circumvents the need to directly constrain the latent representation distribution of Bongard-Logo images or rigidly model this distribution into predefined formats. Experimental results validate the rationality and effectiveness of the PMoC design approach. Consequently, PMoC, enhanced by Pose-Transformer, has exhibited its advantage in tackling the Bongard-Logo problem.

### v8

@@ IV-B Straw Pose-Tansformer
Pose-Transformer, during deployment, exerts considerable pressure on computational resources. The establishment and computation of the mapping matrices, denoted as $\{W_{ijk}|i\in[1,n],j\in[1,m],k\in[1,n],W_{ijk}\in R^{(d/m)\times d}\}$, incur significant resource overhead. Motivated by this challenge, this paper explores the lightweighting of Pose-Transformer. The crux of Pose-Transformer lies in computing compressed pose vectors and subsequently employing them to modify and refine the logical inputs $\{V_{i}|i\in[1,n],V_{i}\in R^{d}\}$ to some extent. The computation of the logical outputs in Pose-Transformer’s pose matrix mapping follows the methodology outlined below:
@@
It is worth noting that the placeholders marked with “_” in the formula represent numbering for on-demand broadcasting. Multiple levels of summation and mapping can incur significant computational overhead. By designing parameter sharing, we aim to eliminate partial logical summation and mapping, thereby achieving the purpose of reducing computational cost. From a logical hierarchy perspective, only the logical mapping at the $i$-th level possesses the potential for indentation, whereas the other logical levels exhibit stringent mapping requirements. In other words, this paper posits that the pose matrices between equivalent local vectors across different tokens merit weight sharing. The planned target expression of pose matrix mapping after indentation is:
@@
This paper aims to simplify the Pose-Transformer in a manner consistent with the aforementioned expression. We treat the mapping matrix $W_{ijk}$ as a product of the mapping matrix $\tilde{W}_{jk}$ multiplied by $i$ different coefficients, specifically expressed as $W_{ijk}\tilde{=}w_{i}\cdot\tilde{W}_{\_jk}$. This transformation significantly reduces the number of mapping matrices involved. We argue that it is more reasonable to combine $w_{i}$ with $H_{ij\_}$ and treat them as a weighted sum of $H_{ij\_}$, rather than treating $w_{i}$ as a set of optimizable preliminary coefficients for mapping matrices. Furthermore, treating this process as computing an attention-based weighted summation for $H_{ij\_}$ is even more elegant and effective. This solution aligns well with the formula presented:
@@
Notably, the aforementioned approach necessitates additionally computing attention-based weighted summation between the standard Transformer-Encoder and pose mapping. In other words, between equations (5) and (6) lies the necessity to incorporate an additional computational step, which prolongs the training time, contradicting the intended purpose. This paper utilizes a Transformer-Encoder with an additional mask mechanism to integrate the Transformer encoding process with the additional attention-based weighted summation process.
Specifically, this paper introduces an optimizable vector, denoted as $\tilde{V}_{0}\in R^{d}$, for the embedding of logical inputs $\{\tilde{V}_{i}|i\in[1,n]\}$. After the embedded logical inputs $\{\tilde{V}_{i}|i\in[0,n]\}$ propagate through a single Transformer-Encoder layer, while applying a mask to the self-attention weights at the position corresponding to $\tilde{V}_{0}$ to prevent its value from contributing to the attention results of the Transformer-Encoder, they produce an updated set of vectors $\{V_{i}|i\in[0,n]\}$ representing the attention-weighted outputs. This process effectively entails the concurrent computation of two procedures: the self-attention calculation within $\{\tilde{V}_{i}|i\in[1,n]\}$ and the cross-attention computation between $\tilde{V}_{0}$ and $\{\tilde{V}_{i}|i\in[1,n]\}$. Accordingly, $\{V_{i}|i\in[1,n]\}$ represents the self-attention outcome of $\{\tilde{V}_{i}|i\in[1,n]\}$, while $V_{0}$ corresponds to the weighted sum of the vectors within $\{\tilde{V}_{i}|i\in[1,n]\}$, where the weights are determined by their attention similarity scores with $\tilde{V}_{0}$. This masked attention can be expressed as Figure 15. In Figure 15, we show a simplified case with only three vectors in $\{\tilde{V}_{i}\}$ for clarity.
Figure 15: Masked attention
We replace the tedious sequential process with this masked attention mechanism, thereby streamlining the computational workflow.
Due to the inherent self-attention mechanism and masking, the vector $\tilde{V}_{0}$ assimilates information from the logical inputs $\{\tilde{V}_{i}|i\in[1,n]\}$, transforming into global vector $V_{0}$. Within this global vector, we establish a collection of mapping matrices $\{\tilde{W}_{jk}|j\in[1,m],k\in[1,n],\tilde{W}_{jk}\in R^{(d/m)\times d}\}$ for each attention head $\{H_{0j}|j\in[1,m],H_{0j}\in R^{(d/m)}\}$ in it.
@@
The pose vectors $\{p_{jk}|j\in[1,m],k\in[1,n],p_{jk}\in R^{d}\}$ is achieved by multiplying the attention heads with the mapping matrices. These pose vectors, after summation and squashing, are added to the vector set $\{V_{i}|i\in[1,n],V_{i}\in R^{d}\}$ to obtain the output of new pose matrix mapping block $\{O_{i}|i\in[1,n],O_{i}\in R^{d}\}$. This process outlines the computation process of the pose embedding matrix block within the “Straw-Pose-Transformer” architecture. The entire calculation process of new “pose matrix embedding” can be expressed as fallows:
@@
In the formula, the process of the Masked Transformer-Encoder to obtain the output is denoted as “Masked T-Encoder”. In the Straw-Pose-Transformer architecture, the outputs from each pose matrix embedding block, denoted as $\{O_{k}|k\in[1,n]\}$, undergo a concatenation process with the global vector $V_{0}$. This modified set $\{V_{0},O_{k}|k\in[1,n]\}$, now represented as $\{\tilde{V}_{i}|i\in[0,n]\}$ , is then propagated into the subsequent logical layer for further processing. However, if the subsequent logical layer is the final layer, the global vector $V_{0}$ is no longer appended.
In this way, we can reconstruct the pose matrix embedding process according to the formula (10). The feedforward process of new pose matrix embedding is illustrated in the figure 16. And figure 17 illustrate the structure of the Straw-Pose-Transformer.
Figure 16: The feedforward process of a new pose matrix embedding block
@@
In addressing the Bongard-Logo problem, this paper introduces a novel baseline, PMoC. To demonstrate the adaptability of the Straw-Pose-Transformer in solving the Bongard-Logo problem, the Transformer-Encoder component within PMoC is replaced with the Straw-Pose-Transformer. Furthermore, given the remarkable performance achieved by the predecessor RS-Tran[30] in solving the RPM problem, this paper does not design an additional baseline network to prove the superiority of the Pose-Transformer in addressing the RPM problem. This paper aims to demonstrate the advantage of Pose-Transformer over the standard Transformer-Encoder for RPM problems by achieving improved accuracy through the simple replacement of the Transformer-Encoder in RS-Tran with Pose-Transformer.
The RS-Tran[30] primarily consists of Transformer-Encoder structures, and in this study, we replace it with the Pose-Transformer and Straw-Pose-Transformer and conduct experiments. The modified RS-Tran is renamed as RS-Pose-Tran and RS-Straw-Pose-Tran. Due to the fundamental modifications made to the RS-Tran architecture, RS-Pose-Tran and RS-Straw-Pose-Tran can be regarded as models designed based on our novel structure, incorporating inductive biases inherent in the RPM problem[41].
@@ VI-A Experiment on Bongard-Logo
All our experiments on Bongard-Logo are implemented in Python using the PyTorch[42] framework. In the experiments conducted for PMoC and SBSD, the Adam[43] optimizer was selected with a learning rate set to 0.001, a learning rate decay of 0.995, a weight decay of 0.0001, and a batch size of 40 for PMoC and SBSD. Specifically, in the experiment combining PMoC with Straw-Pose-Transformer, the batch size was set to 20.
We conducted experiments on the Bongard-logo dataset, utilizing SBSD, PMoC, and the integration of PMoC with Straw-Pose-Transformer. The experimental results are documented in Table I. Due to the propensity of the serial convolutional scanning mechanism and attention mechanism to experience attention collapse when combined, we adopted the parameters of the SBDC model as pre-trained parameters for the PMoC convolutional layer. Notably, under the experimental conditions of this study, which involved the use of four A100 Graphics Processing Units (GPUs), the integration of PMoC with Pose-Transformer posed significant deployment challenges. This served as one of the key reasons and motivations behind our decision to develop a lightweight approach for the Pose-Transformer in this paper.
@@
This paper argues that the adopted data augmentation methods do not explicitly alter the core reasoning difficulty or logical complexity of the Bongard-Logo problem. These augmentation techniques, from an intuitive perspective, neither add nor remove any preexisting concepts. Even if these data augmentation methods implicitly introduce additional and unknown concepts to the Bongard-Logo dataset, it is important to note that such concepts are not present in the test set, as no data augmentation has been applied to it. Furthermore, they do not provide guidance for generalization issues such as NV and CM, as the solutions to these problems are not directly reflected in the HD problem after applying such augmentations. Additionally, the data augmentation approach employed in this study is uniformly applied to all 14 images within each Bongard-Logo instance, maintaining the nature of few-shot learning (i.e., without increasing the number of images within an instance) and without introducing additional priors (such as rotational and flipping invariance of individual images within Bongard-Logo instances). These augmentation methods indeed enrich the expressions of concepts at the pixel level, which is precisely the challenge we anticipate for PMoC through data augmentation and expanded data volume. Enabling multiple expressions for concepts is precisely the original intention behind the establishment of the Bongard-Logo dataset[16].
@@ VI-B Experiment on RPM
Given that I-RAVEN[19] represents a more advanced and rigorous benchmark compared to RAVEN[17], the experiments in this paper were solely conducted on the I-RAVEN dataset. We conducted experiments on I-RAVEN and PGM [18] by sequentially replacing the Transformer-Encoder structure within RS-Tran [30] with Pose-Transformer and Straw-Pose-Transformer to obtain RS-Pose-Tran and RS-Straw-Pose-Tran. The results of I-RAVEN are recorded in Table III and PGM is recorded in Table IV. In RS-Tran, multiple layers of the standard Transformer-Encoder are stacked. However, considering our belief in the significant performance advantages of (Straw) Pose-Transformer compared to the standard Transformer-Encoder, we only stacked half of the logical layers compared to RS-Tran when creating the combined models RS-Pose-Tran and RS-Straw-Pose-Tran. Additionally, the parameter settings and dataset configurations for experiments related to RS-Pose-Tran were maintained strictly consistent with those of RS-Tran. This setup ensures that the comparison results among the three models are more convincing.
@@
In Table III and IV, we observed that compared to the Transformer encoder, the Pose-Transformer exhibited a greater aptitude for abstract reasoning problems, particularly in addressing RPM and Bongard-Logo challenges. Furthermore, it is noteworthy that we refrained from conducting experiments with RS-Pose-Tran on the PGM dataset. This decision was made due to the fact that, compared to RS-Tran operating on the RAVEN dataset, the computational overhead of RS-Tran doubled when applied to the PGM dataset. However, such a significant increase in computational requirements was prohibitive for RS-Pose-Tran, rendering it unfeasible for deployment on our current equipment and facilities. Consequently, this motivated us to introduce the Straw-Pose-Transformer in this study, a lightweight variant of the Pose-Transformer.
@@
From the perspective of experimental results, it is evident that our lightweight approach, RS-Straw-Pose-Tran, has achieved success without compromising the performance of RS-Pose-Tran to a significant degree. Moreover, the incorporation of the pose mapping matrix into the Transformer-Encoder has proven effective, leading to a substantial enhancement in the performance of RS-Tran. Even without taking into account the improvements relative to RS-Tran, the results achieved by RS-Straw-Pose-Tran are indeed noteworthy and can be classified as remarkable.
@@ VII Conclusion
The characteristics of deep networks as black-box intelligent systems, coupled with their optimization methods based on steepest gradient descent, create a gap between the theory and implementation of deep networks. In this paper, we report the design process of PMoC, which represents a “compromise” and “indirect” approach to neural network architecture design philosophy. By employing a unique probabilistic modeling technique, PMoC cleverly circumvents many of the theoretical and practical discrepancies associated with deep learning. This allows PMoC to offer a probabilistic-model-based solution for visual clustering problems. Its effectiveness has been validated on the renowned clustering reasoning problem: Bongard-Logo. This paper advocates for the adoption of similar indirect strategies and compromise approaches when addressing similar visual clustering problems or tasks requiring the estimation of unknown distributions.
As a bonus, we have designed the Pose-Transformer, a variant of the attention-based Transformer, which has excelled in logical reasoning tasks. Both RS-Tran[30] and the PMoC utilize the Transformer as their core backbone structure, and they have achieved notable successes in abstract reasoning problems. This demonstrates the remarkable logical analysis capability inherent in the Transformer. By replacing the Transformer structure in them with our Pose-Transformer, we have observed a direct and significant improvement in their accuracy. This paper encourages the consideration of our versatile hot-swappable structure when addressing visual abstract reasoning problems.